Use this new algebraic method to find equations for lines with the following properties:
A. A slope of -3, passing through the point (15,-50)
B. A slope of 0.5 with an x-intercept of (28,0)

Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

Presumably, one wants the point-slope formula
y-y1=m(x-x1)
For the first
y-(-50)=-3(x-15)
y+50=-3x+45
y=-3x-5; -50=-45-5, check
--------------------------
y=0.5(x-28)
y=(1/2)x-14; 28(1/2)-14=0, check

Answer by Alan3354(69443)

What "new algebraic method?"